OpenStudy (anonymous):

In a bolt factory, machines A, B, and C manufacture 25%, 35% and 40% of the total output, respectively. Of their total of the output, 5%, 4% and 2% respectively are defective bolts. A bolt is drawn at random and is found to be defective. What is the probability that it is manufactured by machine B?

OpenStudy (anonymous):

4/35

OpenStudy (kropot72):

 \[P(A\ defective)=0.25\times0.05\] \[P(B\ defective)=0.35\times0.04\] \[P(C\ defective)=0.4\times0.12\] The probability that the defective bolt is made by machine B is given by: \[\large P(made\ by\ B)=\frac{P(B\ defective)}{P(A\ defective)+P(B\ defective)+P(C\ defective)}\]

OpenStudy (anonymous):

140/100

OpenStudy (anonymous):

0.14/0.2005

OpenStudy (kropot72):

Not really. Lets look at the numerator value: \[\large P(B\ defective)=0.35\times0.04=what\ is\ your\ result?\]

OpenStudy (anonymous):

0.014/0.0745

OpenStudy (anonymous):

0.014

OpenStudy (kropot72):

The numerator is correct now. What are you results for: P(A defective) = ? P(C defective) = ?

OpenStudy (anonymous):

0.0125,0.0008

OpenStudy (anonymous):

P(C defective)=0.4×0.12 u wrote there instead of 0.02.thats why i got confused

OpenStudy (anonymous):

0.014/0.0345

OpenStudy (kropot72):

P(A defective) = 0.0125 P(B defective) = 0.0140 P(C defective) = 0.0080 Sorry for typo earlier. 0.014/0.0345 is the correct fraction. Now you can finally calculate the result.

OpenStudy (anonymous):

so 28/69

OpenStudy (anonymous):

why we are multiplying the output with defective.i thought defective is given directly

OpenStudy (anonymous):

0.35*0.04 why?.do u understand my question

OpenStudy (anonymous):

4/35 itself show it is manufactured by b

OpenStudy (kropot72):

Let event B be that a bolt came from machine B. Let event D be the bolt is defective. P(B) = 0.35 P(D) = 0.04 The probability of B and D is: \[P(B \cap D)=P(B) \times P(D)=0.35\times0.04\]

OpenStudy (anonymous):

nice:-)

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

when we use this formula.

OpenStudy (kropot72):

The formula applies when the two events are independent.